# Knowledge Distillation with Refined Logits

### v2

@@ Abstract
Recent research on knowledge distillation has increasingly focused on logit distillation because of its simplicity, effectiveness, and versatility in model compression. In this paper, we introduce Refined Logit Distillation (RLD) to address the limitations of current logit distillation methods. Our approach is motivated by the observation that even high-performing teacher models can make incorrect predictions, creating a conflict between the standard distillation loss and the cross-entropy loss. This conflict can undermine the consistency of the student model’s learning objectives. Previous attempts to use labels to empirically correct teacher predictions may undermine the class correlation. In contrast, our RLD employs labeling information to dynamically refine teacher logits. In this way, our method can effectively eliminate misleading information from the teacher while preserving crucial class correlations, thus enhancing the value and efficiency of distilled knowledge. Experimental results on CIFAR-100 and ImageNet demonstrate its superiority over existing methods. The code is provided at https://github.com/zju-SWJ/RLD.
@@ 1 Introduction
Knowledge distillation (Gou et al. 2021) utilizes pre-trained high-performing teacher models to facilitate the training of a more compact student model. Compared to other model compression methods, such as pruning and quantization (Choudhary et al. 2020), knowledge distillation exhibits fewer constraints on the model architecture. This flexibility significantly broadens its applicability, contributing to its increasing prominence in recent research.
@@
Hinton et al. (2015) were the first to introduce the concept of logit distillation. It is designed to align the logits of teacher and student models following the softmax operations, using Kullback-Leibler (KL) divergence as the alignment measure. Most subsequent research has maintained the original concept of logit distillation, instead focusing on exploring feature distillation (Romero et al. 2014; Komodakis and Zagoruyko 2017; Tian, Krishnan, and Isola 2020; Chen et al. 2021, 2022; Wang et al. 2022) in more depth, which specifically investigates the process of selecting and aligning intermediate-level features between teacher and student models. However, the potential architectural disparity between teacher and student models poses a significant challenge for feature alignment. This is mainly due to the fact that different architectures extract different features (Wang et al. 2022). Moreover, the extensive diversity in feature selection further amplifies the complexity of feature distillation and leads to an increase in training time in distillation (Chen et al. 2022). Recently, by decoupling the classical logit distillation loss, the study by Zhao et al. (2022) demonstrates that logit distillation can yield results that are on par with, or even superior to, those of feature distillation. Consequently, logit distillation has garnered considerable attention in the research community, thanks to its simplicity, effectiveness, and versatility.
@@
However, despite the impressive achievements of recent logit distillation approaches (Li et al. 2023; Jin, Wang, and Lin 2023; Sun et al. 2024a), most overlook the impact of teacher prediction accuracy on the training process. Specifically, incorrect teacher predictions lead to a conflict between teacher loss and label loss, which may severely impede the potential for further enhancements of the student models. Existing correction-based distillation approaches (Wen, Lai, and Qian 2021; Cao et al. 2023; Lan et al. 2024) consistently modify the teacher logits (target) using label information. They either exchange the values between the predicted maximum class and the true class (Wen, Lai, and Qian 2021) (the swap operation) or amplify the proportion of the true class within the predicted probabilities (Cao et al. 2023; Lan et al. 2024) (the augment operation). We argue that such approaches may alter the correlation among classes, as exemplified in Figure 1. This disruption can obstruct the transmission of “dark knowledge” (Hinton, Vinyals, and Dean 2015) and hinder performance improvements.
@@
Figure 1: A toy example of existing correction-based distillation approaches. Classes represented by the same color are closely related. The image displayed is a “lion”, yet the teacher model incorrectly classifies it as the “forest”. Both the swap and augment operations disrupt the close correlation between “lion” and “tiger”. A more detailed example of class correlation is provided in the appendix.
@@
We introduce Refined Logit Distillation (RLD) to address these challenges. RLD aims to alleviate the loss conflicts that arise in student training by empowering student models to assimilate valuable knowledge from their teachers. Specifically, RLD consists of two types of knowledge, sample confidence (SC) and masked correlation (MC). Sample confidence refers to the binary probabilities derived from logits. In the context of the teacher model, SC comes from the probability associated with the predicted class and the probabilities of the remaining classes. This metric encapsulates the teacher’s prediction confidence for the current sample and can be employed to guide the student model. Considering the possible inaccuracies in the teacher’s prediction, we align the student’s true class prediction probability with that of the teacher’s. This alignment not only mitigates the mistakes by the teacher, but also guides the student model toward achieving a comparable level of confidence for the current sample. Moreover, it can effectively prevent over-fitting. Masked correlation denotes our dynamic approach for selecting a subset of classes for teacher-student alignment. It is designed to mitigate the influence of potentially incorrect teacher predictions on student models while conveying essential class correlation. More specifically, MC involves masking all classes within the teacher logits that have equal or superior rankings compared to the true class. In essence, fewer classes are used for distillation when the teacher makes more mistakes (i.e., lower ranking for the true class), and more classes are used when it makes fewer mistakes. Using these two complementary types of refined knowledge, the student can achieve better performance.
Our contributions can be summarized as follows:
@@
  • We reveal that prevalent distillation approaches fail to account for the effects of incorrect teacher predictions, and existing correction-based strategies tend to ruin the valuable class correlation.
@@
  • We introduce a novel logit distillation approach termed Refined Logit Distillation (RLD) to prevent over-fitting and mitigate the influence of incorrect teacher knowledge, while preserving the essential class correlation.
@@
Figure 2: An overview of our proposed Refined Logit Distillation (RLD). In RLD, the teacher model imparts two types of knowledge, denoted as “sample confidence” and “masked correlation”, to the student model. Both kinds of knowledge are obtained from logits, thus the distillation process does not introduce intermediate layer features.
@@ 2 Related Work
The application of knowledge distillation historically concentrated on the image classification task, and progressively extended to a wider range of tasks, including semantic segmentation (Liu et al. 2020; Shu et al. 2021; Yang et al. 2022; Sun et al. 2023b) and image generation (Salimans and Ho 2022; Sun et al. 2023a; Meng et al. 2023) within the realm of computer vision. Traditional knowledge distillation typically involves a single teacher and a single student model. As the field evolves, a variety of other paradigms have been proposed, such as online distillation (Chen et al. 2020; Wu and Gong 2021), multi-teacher distillation (Yuan et al. 2021; Zhang, Chen, and Wang 2022), and self-distillation (Furlanello et al. 2018; Kim et al. 2021; Sun et al. 2024b). Since traditional knowledge distillation remains the core foundation of research in this area, we will focus solely on such methods in the discussion below.
In image classification task, existing algorithms can be broadly classified into three categories: logit distillation, feature distillation (Komodakis and Zagoruyko 2017; Kim, Park, and Kwak 2018; Heo et al. 2019b; Tian, Krishnan, and Isola 2020; Chen et al. 2021; Wang et al. 2022; Chen et al. 2022), and relation distillation (Park et al. 2019; Liu et al. 2019; Peng et al. 2019). Logit distillation has become the main focus of current research because of its straightforwardness, effectiveness, and adaptability. The initial logit distillation (Hinton, Vinyals, and Dean 2015) leverages KL divergence to align the softened output logits of the teacher and student models, thereby significantly enhancing the performance of the student models. DKD (Zhao et al. 2022) revitalizes logit distillation by decoupling this classical loss, enabling it to perform comparably to feature distillation. MLKD (Jin, Wang, and Lin 2023) leverages multi-level logit knowledge to further enhance model performance. CTKD (Li et al. 2023) introduces the curriculum temperature, applying adversarial training and curriculum learning to dynamically determine the distillation temperature for each sample. LSKD (Sun et al. 2024a) processes the logits to adaptively allocate temperatures between teacher and student and across samples, thereby achieving state-of-the-art performance. However, the effect of incorrect teacher predictions on distillation is rarely considered.
Given that logits are intrinsically related to prediction accuracy, several methods leverage labels to adjust logits prior to the distillation process. LA (Wen, Lai, and Qian 2021) swaps the values of the true and predicted classes to correct the teacher model’s predictions. RC (Cao et al. 2023) amplifies the predicted value for the true class with the highest predicted probability in the student’s output distribution, thereby aiding the student model in making accurate and confident predictions. LR (Lan et al. 2024) combines one-hot labels with the teacher’s soft labels to produce a new, precise target for distillation. However, as previously demonstrated, these methods may disrupt class correlations, which can hinder performance improvement.
@@ 3 Preliminaries
By combining Eqs. 2 and 3, we get the classical logit distillation loss for stochastic gradient descent. Such an approach has been experimentally shown to perform better than training solely with labels.
@@ 4.1 Sample Confidence Distillation
Sample confidence (SC) represents the binary distribution $b$ derived from the logits. It encapsulates the model confidence for each sample, thereby aiding the student model in generating high-confidence predictions for the true class, without unduly restricting the distribution for other classes.
In the context of teacher knowledge, one component of the SC is the maximum predicted probability value $\hat{p}_{\text{max}}^{\text{T}}$, while the other is the sum of predicted probabilities for the remaining classes. However, the student SC consists of the predicted probability for the true class $\hat{p}_{\text{true}}^{\text{S}}$, and the other includes the predicted probabilities for the remaining classes. They can be summarized in the following formulas:
@@
Figure 3: Toy examples elucidating the variances across methods when the teacher and student models are well-aligned. (a) Since entropy does not preserve the probability rank of different classes, it may cause incorrect student prediction. (b) Our proposed binary sample confidence guarantees a high probability for the true class and eliminates the necessity for the student model to match the intact class probability distribution to that of the teacher model. (c) Masked correlation imposes fewer constraints on the student prediction than traditional knowledge distillation. For the classes that are masked, their probabilities can diverge entirely from those of the teacher model.
Although entropy could potentially be employed to measure sample confidence, it would not be appropriate for our purpose. The aim of SC is to enable the student model to maintain a similar level of confidence for the true class without hindering the predicted probabilities of the other classes. However, a straightforward transfer of entropy does not impose this constraint on the true class of the student model, as shown in Figure 3(a) and (b).
@@ 4.2 Masked Correlation Distillation
$L_{\text{SCD}}$ Mask Accuracy
$M_{\text{g}}$ $M_{\text{ge}}$
75.50
75.64
75.53
76.64
Table 1: Top-1 accuracy (%) on the CIFAR-100 validation set when training with different masking strategies. The teacher is ResNet32$\times$4, and the student is ResNet8$\times$4.
Masked correlation (MC) denotes the probability distribution acquired after dynamically masking certain classes. As shown in Figure 3(c), this masking operation relieves the student model from aligning incorrect class rankings, thereby allowing the student model to generate very different output from the teacher without incurring a large loss. Moreover, preserving partial class probabilities empowers the student model to learn valuable class correlation, consequently enhancing the model’s performance.
@@
Although masking off all classes that have values greater than (denoted as “g”) the true class would similarly remove the misinformation and preserve the class correlation, this masking strategy (denoted as $M_{\text{g}}$) inadvertently incorporate true class-related knowledge into both masked correlation and sample confidence. As demonstrated in the Table 1, this could potentially introduce conflicts among losses and impact the model performance. Hence, we opt for $M_{\text{ge}}$ as the masking strategy.
@@
When the teacher model makes a more accurate prediction (ranking the true class higher), only a few classes are subjected to the masking operation. It allows the majority of class correlation to be preserved and transferred to the student model. Conversely, if the teacher’s prediction is less accurate, the majority of classes are masked. As a result, the student model learns less knowledge, thereby reducing the potential for misinformation to mislead the training process. It also gives the student model more freedom to make predictions for masked classes that differ significantly from those of the teacher model.
@@ 4.3 Refined Logit Distillation
By combining Eqs. 26, and 9, we obtain the final loss for RLD, which is
@@
where $\alpha$ and $\beta$ are hyper-parameters to adjust the importance of sample confidence and masked correlation, respectively. A detailed algorithm is shown in Algorithm 1, where temperature $\tau$ is omitted for concise.
Algorithm 1 Refined Logit Distillation
Input: Training dataset $D=\{\left(x_{n},y_{n}\right)\}^{N}_{n=1}$; Pre-trained teacher model $\theta^{\text{T}}$; Loss hyper-parameters $\alpha$ and $\beta$
Output: Student model $\theta^{\text{S}}$
1:  Randomly initialize $\theta^{\text{S}}$.
2:  while not converged do
3:     for $(x,y)$ in $D$ do
4:        Propagate $x$ forward through $\theta$ to get logits $z$.
5:        Calculate $L_{\text{CE}}$ using $z^{\text{S}}$ and $y$ (Eqs. 1 and 2).
6:        Obtain sample confidence $b$ using $z$ (Eqs. 4 and 5).
7:        Calculate $L_{\text{SCD}}$ using $b^{\text{T}}$ and $b^{\text{S}}$ (Eq. 6).
8:        Obtain mask $M_{\text{ge}}$ using $z^{\text{T}}$ and $y$ (Eq. 7).
9:        Obtain correlation $\tilde{p}$ using $z$ and $M_{\text{ge}}$ (Eq. 8).
10:        Calculate $L_{\text{MCD}}$ using $\tilde{p}^{\text{T}}$ and $\tilde{p}^{\text{S}}$ (Eq. 9).
11:        Calculate $L_{\text{RLD}}$ with $\alpha$ and $\beta$ (Eq. 10).
12:        Backward $L_{\text{RLD}}$ and update $\theta^{\text{S}}$.
13:     end for
14:  end while
15:  return $\theta^{\text{S}}$
Next, we elucidate the role of each loss function, which collaboratively enhances the model performance.
  • $L_{\text{CE}}$ encourages the student model to generate the highest probability for the true class, but it may cause over-fitting when deployed independently.
  • $L_{\text{SCD}}$ expects the student model to attain a reasonable level of confidence for the true class, thus averting over-fitting. However, when used in isolation, it could yield the probability assigned to one of the remaining classes exceeds that of the true class. When $L_{\text{CE}}$ is amalgamated with $L_{\text{SCD}}$, the true class manages to retain the highest and most suitable probability. However, this combination falls short in transferring knowledge about the remaining classes.
  • $L_{\text{MCD}}$ lacks information about the true class (owing to the consistent masking of the true class), but it has the ability to eliminate misinformation and transfer valuable class correlation to the student model. By integrating $L_{\text{CE}}$ and $L_{\text{SCD}}$ with $L_{\text{MCD}}$, we can ensure the delivery of substantial valuable knowledge.
@@
Intriguingly, although our approach and DKD (Zhao et al. 2022) consider logit distillation from distinct perspectives, they become equivalent when the teacher model consistently makes accurate predictions. Besides, DKD does not explicitly elucidate why it is effective to underscore the non-target class knowledge, while our approach implies that by masking the true class during distribution alignment, the student model is provided with greater autonomy to adjust the ranking of the true class, thereby facilitating more accurate predictions.

### v3

@@ Abstract
Recent research on knowledge distillation has increasingly focused on logit distillation because of its simplicity, effectiveness, and versatility in model compression. In this paper, we introduce Refined Logit Distillation (RLD) to address the limitations of current logit distillation methods. Our approach is motivated by the observation that even high-performing teacher models can make incorrect predictions, creating an exacerbated divergence between the standard distillation loss and the cross-entropy loss, which can undermine the consistency of the student model’s learning objectives. Previous attempts to use labels to empirically correct teacher predictions may undermine the class correlations. In contrast, our RLD employs labeling information to dynamically refine teacher logits. In this way, our method can effectively eliminate misleading information from the teacher while preserving crucial class correlations, thus enhancing the value and efficiency of distilled knowledge. Experimental results on CIFAR-100 and ImageNet demonstrate its superiority over existing methods. Our code is available at https://github.com/zju-SWJ/RLD.
@@ 1 Introduction
Knowledge distillation utilizes pre-trained high-performing teacher models to facilitate the training of a compact student model [12]. Compared to other model compression methods, such as pruning and quantization [5], knowledge distillation exhibits fewer constraints on the model architecture. This flexibility significantly broadens its applicability, contributing to its increasing prominence in recent research.
@@
Hinton et al[12] were the first to introduce the concept of logit distillation. It is designed to align the logits of teacher and student models following the softmax operations with the Kullback-Leibler (KL) divergence. Most subsequent research has maintained the original concept of logit distillation, instead focusing on exploring feature distillation [29, 18, 39, 4, 3, 40] in more depth by selecting and aligning intermediate-level features between teacher and student models. However, the potential architectural disparity between teacher and student models poses a significant challenge for feature alignment. This is mainly due to the fact that different architectures extract different features [40]. Moreover, the extensive diversity in feature selection further amplifies the complexity of feature distillation and leads to an increase in training time in distillation [3]. Recently, by decoupling the classical logit distillation loss, Zhao et al[50] demonstrate that logit distillation can yield results that are on par with, or even superior to, those of feature distillation. Consequently, logit distillation garnered considerable attention in the research community, thanks to its simplicity, effectiveness, and versatility.
@@
Despite the impressive achievements, most of the recent logit distillation approaches [21, 15, 35] overlook the impact of teacher prediction correctness on the training process. Specifically, incorrect teacher predictions lead to an exacerbated divergence between teacher loss and label loss, which may severely impede the potential enhancements of the student models. Existing correction-based distillation approaches [42, 1, 20] consistently modify the teacher logits (target) using label information. They either exchange the values between the predicted maximum class and the true class [42] (the swap operation) or amplify the proportion of the true class within the predicted probabilities [1, 20] (the augment operation). We argue that such approaches may alter the correlations among classes, as exemplified in Figure 1. This disruption can obstruct the transmission of “dark knowledge” [12] and hinder performance improvements.
@@
Figure 1: A toy example of existing correction-based distillation approaches. Classes represented by the same color are highly correlated and should be ranked closely. The image displayed is a “lion”, yet the teacher model incorrectly classifies it as the “forest”. Both the swap and augment operations disrupt the close correlation between “lion” and “tiger”. A more detailed example of class correlation is provided in the Appendix.
@@
In this paper, we introduce Refined Logit Distillation (RLD) to address these challenges. In classification tasks, the true class probability dictates prediction correctness, while class correlations capture high-level semantic relationships that influence classification tendencies. Accordingly, RLD consists of two types of knowledge, sample confidence (SC) and masked correlation (MC). Sample confidence refers to the binary probabilities derived from logits. As for the teacher model, SC comes from the probability associated with the predicted class and the probabilities of the remaining classes. It encapsulates the teacher’s prediction confidence for the current sample and is employed to guide the student model. Considering the possible inaccuracies in the teacher’s prediction, we align the student’s true class probability with teacher’s predicted class probability. This alignment not only mitigates the teacher’s mistakes, but also guides the student model toward achieving a comparable level of confidence for the current sample. Moreover, it effectively prevents over-fitting. Masked correlation denotes our dynamic approach for selecting a subset of classes for teacher-student alignment. It is designed to mitigate the influence of potentially incorrect teacher predictions on student models while conveying essential class correlations. More specifically, MC involves masking all classes within the teacher logits that have equal or superior rankings compared to the true class. In essence, fewer classes are used for distillation when the teacher makes more mistakes, and more classes are used when it makes fewer mistakes. Using these two complementary types of refined knowledge, the student can achieve better performance.
Our contributions are summarized as follows:
@@
  • We reveal that prevalent distillation approaches fail to account for the effects of incorrect teacher predictions, and existing correction-based strategies tend to ruin the valuable class correlations.
@@
  • We introduce a novel logit distillation approach termed Refined Logit Distillation (RLD) to prevent over-fitting and mitigate the influence of incorrect teacher knowledge, while preserving the essential class correlations.
@@
Figure 2: An overview of our proposed Refined Logit Distillation (RLD). In RLD, the teacher model imparts two types of knowledge, denoted as “sample confidence” and “masked correlation”, to the student model. The binary sample confidence encapsulates the model confidence for each sample, which helps the student model generate proper-confidence predictions for the true class. The masked correlation denotes the probability distribution acquired after dynamically masking certain classes, which helps to remove misleading information and preserve valuable class correlations during distillation. Both kinds of knowledge are obtained from logits, thus the distillation process does not introduce intermediate layer features.
@@ 2 Related Work
The application of knowledge distillation historically concentrated on the image classification task, and progressively extended to a wider range of tasks, including semantic segmentation [23, 33, 44, 37] and image generation [31, 25, 36] within the realm of computer vision. Traditional knowledge distillation typically involves a single teacher and a single student model. As the field evolves, a variety of other paradigms have been proposed, such as online distillation [2, 43], multi-teacher distillation [45, 48], and self-distillation [7, 17]. Since traditional knowledge distillation remains the core foundation of research in this area, we will focus solely on such methods in the discussion below.
In image classification task, existing algorithms can be broadly classified into three categories: logit distillation [12, 50, 15, 21, 35], feature distillation [18, 16, 11, 39, 4, 40, 3], and relation distillation [26, 22, 28]. Logit distillation has become the main focus of current research because of its straightforwardness, effectiveness, and adaptability. The initial logit distillation [12] leverages KL divergence to align the softened output logits of the teacher and student models, thereby significantly enhancing the performance of the student models. DKD [50] revitalizes logit distillation by decoupling this classical loss, enabling it to perform comparably to feature distillation. MLKD [15] leverages multi-level logit knowledge to further enhance model performance. CTKD [21] introduces the curriculum temperature, applying adversarial training and curriculum learning to dynamically determine the distillation temperature for each sample. LSKD [35] processes the logits to adaptively allocate temperatures between teacher and student and across samples, thereby achieving state-of-the-art performance. However, the effect of incorrect teacher predictions on distillation is rarely considered.
Given that logits are intrinsically related to prediction correctness, several methods leverage labels to adjust logits prior to the distillation process. LA [42] swaps the values of the true and predicted classes to correct the teacher model’s predictions. RC [1] adds the maximum value in the student’s output to the true class, thereby aiding the student model in making accurate and confident predictions. LR [20] combines one-hot labels with the teacher’s soft labels to produce a new, precise target for distillation. However, as previously demonstrated in Figure 1, these methods may disrupt class correlations, which can hinder performance improvement.
@@ 3 Preliminaries
By combining Equations 2 and 3, we get the classical logit distillation loss for stochastic gradient descent. Such an approach has been experimentally shown to perform better than training solely with labels.
@@ 4.1 Sample Confidence Distillation
Sample confidence (SC) represents the binary distribution $b$ derived from the logits. It encapsulates the model confidence for each sample, thereby aiding the student model in generating proper-confidence predictions for the true class, without unduly restricting the distribution for other classes.
In the context of teacher knowledge, one component of the SC is the maximum predicted probability value $\hat{p}_{\text{max}}^{\text{T}}$, while the other component is the sum of the predicted probabilities for the remaining classes. In contrast, the student SC consists of two components: the predicted probability for the true class $\hat{p}_{\text{true}}^{\text{S}}$, and the sum of the predicted probabilities for the remaining classes. They can be summarized in the following formulas:
@@
Figure 3(a) more vividly illustrates the aligned knowledge of the teacher and the student when using SCD.
Figure 3: Toy examples elucidating the variances across methods when the corresponding losses are minimal. (a) Sample confidence guarantees a proper probability for the true class and eliminates the necessity for the student model to match the intact class probability distribution to that of the teacher model. (b) Masked correlation imposes fewer constraints on the student prediction than traditional knowledge distillation. For the classes that are masked, their probabilities can diverge entirely from those of the teacher model.
While both $L_{\text{CE}}$ and $L_{\text{SCD}}$ operate on the true class, gradient analysis (temperature $\tau$ omitted) shows that their effects are not entirely identical:
$\begin{gathered}\frac{\partial L_{\mathrm{CE}}}{\partial z_{i}}=p^{\text{S}}_{i}-y_{i},\\ \frac{\partial L_{\mathrm{SCD}}}{\partial z_{i}}=\begin{cases}p^{\text{S}}_{i}-p^{\text{T}}_{\text{max}},&\mbox{$i=\text{true}$},\\ \frac{p^{\text{S}}_{i}(p^{\text{S}}_{\text{true}}-p^{\text{T}}_{\text{max}})}{p^{\text{S}}_{\text{true}}-1},&\mbox{$i\neq\text{true}$}.\end{cases}\end{gathered}$ (7)
Consequently, RLD incorporates both, thereby endowing the distillation process with greater flexibility.
@@ 4.2 Masked Correlation Distillation
Masked correlation (MC) denotes the probability distribution acquired after dynamically masking certain classes. As shown in Figure 3(b), this masking operation relieves the student model from aligning incorrect class rankings, thereby allowing the student model to generate very different output from the teacher without incurring a large loss. Moreover, preserving partial class probabilities empowers the student model to learn valuable class correlations, consequently enhancing the model’s performance.
@@
When the teacher model makes a more accurate prediction (ranking the true class higher), only a few classes are subjected to the masking operation. It allows the majority of class correlations to be preserved and transferred to the student model. Conversely, if the teacher’s prediction is less accurate, the majority of classes are masked. As a result, the student model learns less knowledge, thereby reducing the potential for misinformation to mislead the training process. It also gives the student model more freedom to make predictions for masked classes that differ significantly from those of the teacher model.
@@ 4.3 Refined Logit Distillation
By combining Equations 2, 6 and 10, we obtain the final loss for RLD, which is:
@@
where hyper-parameters $\alpha$ and $\beta$ adjust the importance of sample confidence and masked correlation, respectively.
Type Teacher ResNet32$\times$4 VGG13 WRN-40-2 ResNet56 ResNet110 ResNet110
79.42 74.64 75.61 72.34 74.31 74.31
Student ResNet8$\times$4 VGG8 WRN-40-1 ResNet20 ResNet32 ResNet20
72.50 70.36 71.98 69.06 71.14 69.06
Feature FitNet 73.50 71.02 72.24 69.21 71.06 68.99
AT 73.44 71.43 72.77 70.55 72.31 70.65
RKD 71.90 71.48 72.22 69.61 71.82 69.25
CRD 75.51 73.94 74.14 71.16 73.48 71.46
OFD 74.95 73.95 74.33 70.98 73.23 71.29
ReviewKD 75.63 74.84 75.09 71.89 73.89 71.34
SimKD 78.08 74.89 74.53 71.05 73.92 71.06
CAT-KD 76.91 74.65 74.82 71.62 73.62 71.37
Logit KD 73.33 72.98 73.54 70.66 73.08 70.67
CTKD 73.39 73.52 73.93 71.19 73.52 70.99
DKD 76.32 74.68 74.81 71.97 74.11 71.06
LA 73.46 73.51 73.75 71.24 73.39 70.86
RC 74.68 73.37 74.07 71.63 73.44 71.41
LR 76.06 74.66 74.42 70.74 73.52 70.61
RLD (ours) 76.64 74.93 74.88 72.00 74.02 71.67
Table 1: Top-1 accuracy (%) on the CIFAR-100 validation set when the teacher and student models are homogeneous. The best and second best results of logit distillation are highlighted in bold and underlined text, respectively. For the case where the best result of feature distillation is better than the best result of logit distillation, we highlight it with italic text. The reported results are the mean of three trials.
@@
Although RLD and DKD [50] consider logit distillation from distinct perspectives, they become equivalent when the teacher model consistently makes accurate predictions. Besides, DKD does not explicitly explain why transferring non-target class knowledge (i.e., the probability distribution when the true class is masked, referred to as NCKD) can significantly enhance model performance. Beyond the idea that the class relationships embedded in this knowledge facilitate training, RLD offers a new explanation: when the alignment constraint on target class knowledge (TCKD) is weak, masking the true class during distribution alignment offers the student model greater flexibility to adjust the ranking of the true class. This, in turn, mitigates the negative impact of incorrect teacher knowledge, enabling more accurate predictions.